OpenStudy (anonymous):

What is the slope of the line that passes through the pair of points (2, 9), and (5, 12)? 0 1 3 undefined

OpenStudy (anonymous):

y2-y1/x2-x1

OpenStudy (anonymous):

change in y / change in x (9-12)/(2-5)

OpenStudy (anonymous):

-3/-3 = 1 slope is 1

OpenStudy (anonymous):

medal pls

OpenStudy (anonymous):

(-4-(-7)/(3-5) = 3/-2 = -3/2 -3/2 is final answer -three halves medal and best response pls

OpenStudy (anonymous):

click medal icon right of my name thx ty

OpenStudy (anonymous):

should be --------------------------------------------------------------right here ---------------------------->

OpenStudy (anonymous):

btw am i just giving u answers for ur homework?

OpenStudy (anonymous):

u should feel bad!

OpenStudy (anonymous):

cheater!

OpenStudy (anonymous):

srry ive been having alot of problems for a while

OpenStudy (anonymous):

y-1 = 4(x-5) is answer

OpenStudy (anonymous):

you're* :o

OpenStudy (jojomojo):

hahh, you reported me, no thanks bro

OpenStudy (anonymous):

reported man i dont even know how to use this i thought i gave u a medal wtf :?

OpenStudy (jojomojo):

LOL, well idunno then

OpenStudy (anonymous):

my bad

OpenStudy (jojomojo):

its all good